in three years time Ann will be one third of her mothers age. In another 18 years from now ,Ann will be half her mothers age. how old is her mother now?
Answer by flame8855(424)

A+3 = 1/3 (M+3)
A+18 = (M+18)/2
A+3 = 1/3 (M+3)
-A-18 = -(M+18)/2
-15 = M/3 +1 -M/2 -9
M = 42